The parallel plates of an air filled capacitor with area of area 1.1 x 10^8 m^2 . What must be the width of the plate be if the Capacitance is 1 Farad?
1
Expert's answer
2012-09-27T10:21:24-0400
Using the formula

C = A*e/d

and substituting C = 1 F, A = 1.1 x 10^8 m², e = 8.8 x 10^(-12), we get

d = 9.68*10^(-4) meters.